OpenStudy (hang254):

A pineapple was found to have two sets of opposing spirals formed by its leaves. The clockwise spiral has x leaves. The number of leaves in the counterclockwise direction is 5 more than that in the clockwise direction. These leaf numbers are consecutive numbers in the Fibonacci sequence. Which equation is true for the number of leaves in the two spirals?

OpenStudy (hang254):

Option 1

OpenStudy (hang254):

option 2

OpenStudy (hang254):

option 3

OpenStudy (hang254):

option 4

OpenStudy (hang254):

@ash2326 can you please help me?

OpenStudy (ash2326):

Hey hang, do you know fibonacci sequence?

OpenStudy (hang254):

yes, i do

OpenStudy (hang254):

0, 1, 1, 2, 3, 5, 8, 13, 21, 34, 55, 89, 144, 233, 377, 610, 987, 1597, 2584, 4181, 6765, 10946, 17711, 28657, 46368, 75025, 121393, 196418, 317811

OpenStudy (hang254):

@ash2326 can you please help me?

OpenStudy (ash2326):

Could you write the sequence for me?

OpenStudy (hang254):

You always add the last two numbers to get the next number

OpenStudy (ash2326):

0, 1, 1, 2, 3, 5, 8, 13, 21 notice the terms after 5

OpenStudy (hang254):

what about them?

OpenStudy (ash2326):

The number of leaves in the counterclockwise direction is 5 more than that in the clockwise direction. These leaf numbers are consecutive numbers in the Fibonacci sequence.

OpenStudy (hang254):

ok

OpenStudy (ash2326):

got the idea?

OpenStudy (hang254):

yes, but i dont see how it relates to the problem

OpenStudy (ash2326):

brb

OpenStudy (hang254):

ok

OpenStudy (hang254):

i tried to substitute 13 for x, and none of them work

OpenStudy (hang254):

21, and 34 dont work either

OpenStudy (ash2326):

You know no. of leaves in the counter clock wise is 5 more than the no. in clock wise let x be the number in clock wise x+5 will be the number in anticlockwise

OpenStudy (ash2326):

the hint is in the question, the no. of leaves are consecutive no. in Fibonacci so 0, 1, 1, 2, 3, 5, 8, 13, 21 3 is generated by sum of 2 and 1 so 3=2+1 8 by 5 and 3, so 8=5+3 13 by 8 and 5 , so 13= 8+5 \[\large 8, 13\] does this ring a bell?

OpenStudy (hang254):

yes

OpenStudy (ash2326):

so what's x ?

OpenStudy (hang254):

x=8, 8/5=1.6

OpenStudy (ash2326):

yeah :D did you understand?

OpenStudy (hang254):

yes i did understand.

OpenStudy (hang254):

Thank you for your help :)

OpenStudy (ash2326):

welcome :D